# A Survey on Modern Recommendation System based on Big Data

### 18 Jan 2024 v2

@@ Abstract
This survey provides an exhaustive exploration of the evolution and current state of recommendation systems, which have seen widespread integration in various web applications. It focuses on the advancement of personalized recommendation strategies for online products or services. We categorize recommendation techniques into four primary types: content-based, collaborative filtering-based, knowledge-based, and hybrid-based, each addressing unique scenarios. The survey offers a detailed examination of the historical context and the latest innovative approaches in recommendation systems, particularly those employing big data. Additionally, it identifies and discusses key challenges faced by modern recommendation systems, such as data sparsity, scalability issues, and the need for diversity in recommendations. The survey concludes by highlighting these challenges as potential areas for fruitful future research in the field.
@@ 1 Introduction
In this survey, we examine the escalating popularity and diverse application of recommendation systems in web applications, a topic extensively covered by Zhou et al. [1]. These systems, a specialized category of information filtering systems, are designed to predict user preferences for various items. They play a crucial role in guiding decision-making processes, such as purchasing decisions and music selections, as Wang et al. discuss [2]. A prime example of this application is Amazon’s personalized recommendation engine, which tailors each user’s homepage. Major companies like Amazon, YouTube, and Netflix employ these systems to enhance user experience and generate significant revenue, as noted by Adomavicius et al. and Omura et al. [3, 4]. Figure 1 from Entezari et al. [5] illustrates a modern recommendation system. Additionally, these systems are increasingly relevant in the field of human-computer interaction (HCI), where they enhance interaction efficiency through feedback mechanisms, a topic explored in several studies [6, 7, 8, 9].
Recommendation systems are particularly crucial for certain companies, as their efficiency can lead to substantial revenue generation and competitive advantage, as evidenced in the research by Rismanto et al. and Cui et al. [10, 11]. For instance, Netflix’s “Netflix Prize” challenge aimed to develop a recommender system surpassing their existing algorithm, with a substantial prize to incentivize innovation.
@@
Furthermore, in the domain of big data, recommendation systems are highly prevalent, as detailed by Li et al. [12, 13]. These systems predict user interests in purchasing based on extensive data analysis, including purchase history, ratings, and reviews. There are four widely recognized types of recommendation systems, as identified by Numnonda [14]: content-based, collaborative filtering-based, knowledge-based, and hybrid-based, each with distinct advantages and drawbacks, as Xiao et al. elucidate [15]. For example, collaborative filtering-based systems may face issues such as data sparsity and scalability, as Huang et al. mention [16], and cold-start problems, while content-based systems might struggle to diversify user interests, as noted by Zhang et al. and Benouaret et al. [17, 18].
This paper is organized as follows: Section II provides a comprehensive review of both historical and modern state-of-the-art approaches in recommendation systems, coupled with an in-depth analysis of the latest advancements in the field. Section III discusses the challenges in big data-based recommendation systems, including sparsity, scalability, and diversity, and explores solutions for these challenges. The paper concludes with a summary in Section IV.
@@ 2 Recommendation Systems
Recommendation systems aim to predict users’ preferences for a certain item and provide personalized services [19]. This section will discuss several commonly used recommender methods, such as content-based method, collaborative filtering-based method, knowledge-based method, and hybrid-based method.
@@ 2.1 Content-based Recommendation Systems
The main idea of content-based recommenders is to recommend items based on the similarity between different users or items [20]. This algorithm determines and differentiates the main common attributes of a particular user’s favorite items by analyzing the descriptions of those items. Then, these preferences are stored in this user’s profile. The algorithm then recommends items with a higher degree of similarity with the user’s profile. Besides, content-based recommendation systems can capture the specific interests of the user and can recommend rare items that are of little interest to other users. However, since the feature representations of items are designed manually to a certain extent, this method requires a lot of domain knowledge. In addition, content-based recommendation systems can only recommend based on users’ existing interests, so the ability to expand users’ existing interests is limited.
@@ 2.2 Collaborative Filtering-based Recommendation Systems
Collaborative Filtering-based (CF) methods are primarily used in big data processing platforms due to their parallelization characteristics [21]. The basic principle of the recommendation system based on collaborative filtering is shown in Fig. 2 [22]. CF recommendation systems use the behavior of a group of users to recommend to other users [23]. There are mainly two types of collaborative filtering techniques, which are user-based and item-based.
  • User-based CF: In the user-based CF recommendation system, users will receive recommendations of products that similar users like [24]. Many similarity metrics can calculate the similarity between users or items, such as Constrained Pearson Correlation coefficient (CPC), cosine similarity, adjusted cosine similarity, etc. For example, cosine similarity is a measure of similarity between two vectors. Let $x$ and $y$ denote two vectors, cosine similarity between $x$ and $y$ can be represented by
@@
  • Item-based CF: Item-based CF algorithm predicts user ratings for items based on item similarity. Generally, item-based CF yields better results than user-based CF because user-based CF suffers from sparsity and scalability issues. However, both user-based CF and item-based CF may suffer from cold-start problems [25].
@@
Musto et al. [26]
Volkovs et al.[27]
Mittal et al.[28]
Almaguer et al.[29]
@@
Zhang et al.[30]
Bobadilla et al.[31]
Bobadilla et al.[32]
Rezaimehr et al.[24]
@@
Dong et al.[33]
Gazdar et al.[34]
Alamdari et al.[35]
Cena et al.[36]
@@
Hrnjica et al.[37]
Shokeen et al.[38]
Zagra et al.[39]
Ibrahim et al.[40]
@@ 2.3 Knowledge-based Recommendation Systems
The main idea of knowledge-based recommendation systems is to recommend items to users based on basic knowledge of users, items, and relationships between items [41, 42]. Since knowledge-based recommendation systems do not require user ratings or purchase history, there is no cold start problem for this type of recommendation [43]. Knowledge-based recommendation systems are well suited for complex domains where items are not frequently purchased, such as cars and apartments [44]. But the acquisition of required domain knowledge can become a bottleneck for this recommendation technique [33].
@@ 2.4 Hybrid-based Recommendation Systems
Hybrid-based recommendation systems combine the advantages of multiple recommendation techniques and aim to overcome the potential weaknesses in traditional recommendation systems [45]. There are seven basic hybrid recommendation techniques [40]: weighted, mixed, switching, feature combination, feature augmentation, cascade, and meta-level methods [46, 47]. Among all of these methods, the most commonly used is the combination of the CF recommendation methods with other recommendation methods (such as content-based or knowledge-based) to avoid sparsity, scalability, and cold-start problems [37, 39, 48].
@@ 2.5 Challenges in Modern Recommendation Systems
  • Sparsity. As we know, the usage of recommendation systems is growing rapidly. Many commercial recommendation systems use large datasets, and the user-item matrix used for filtering may be very large and sparse. Therefore, the performance of the recommendation process may be degraded due to the cold start problems caused by data sparsity [49].
  • Scalability. Traditional algorithms will face scalability issues as the number of users and items increases. Assuming there are millions of customers and millions of items, the algorithm’s complexity will be too large. However, recommendation systems must respond to the user’s needs immediately, regardless of the user’s rating history and purchase situation, which requires high scalability. For example, Twitter is a large web company that uses clusters of machines to scale recommendations for its millions of users [38].
  • Diversity. Recommendation systems also need to increase diversity to help users discover new items. Unfortunately, some traditional algorithms may accidentally do the opposite because they always recommend popular and highly-rated items that some specific users love. Therefore, new hybrid methods need to be developed to improve the performance of the recommendation systems [50].
@@ 3 Recommendation System based on Big Data
Big data refers to the massive, high growth rate and diversified information [51, 52]. It requires new processing models to have stronger decision-making and process optimization capabilities [53]. Big data has its unique “4V” characteristics, as shown in Fig. 3 [54]: Volume, Variety, Velocity, and Veracity.
@@ 3.1 Big Data Processing Flow
Big data comes from many sources, and there are many methods to process it [55]. However, the primary processing of big data can be divided into four steps [56]. Besides, Fig. 4 presents the basic flow of big data processing.
@@
  • Data Processing and Integration. The collection terminal itself already has a data repository, but it cannot accurately analyze the data. The received information needs to be pre-processed [57].
  • Data Analysis. In this process, these initial data are always deeply analyzed using cloud computing technology [58].
@@ 3.2 Modern Recommendation Systems based on the Big Data
The shortcomings of traditional recommendation systems mainly focus on insufficient scalability and parallelism [59]. For small-scale recommendation tasks, a single desktop computer is sufficient for data mining goals, and many techniques are designed for this type of problems [60].
@@
However, the rating data is usually so large for medium-scale recommendation systems that it is impossible to load all the data into memory at once [61]. Common solutions are based on parallel computing or collective mining, sampling and aggregating data from different sources, and using parallel computing programming to perform the mining process [62]. The big data processing framework will rely on cluster computers with high-performance computing platforms [63]. At the same time, data mining tasks will be deployed on a large number of computing nodes (i.e., clusters) by running some parallel programming tools [64], such as MapReduce [52, 65]. For example, Fig. 5 is the MapReduce in the Recommendation Systems.
In recent years, various big data platforms have emerged [66]. For example, Hadoop and Spark [52], both developed by the Apache Software Foundation, are widely used open-source frameworks for big data architectures [52, 67]. Each framework contains an extensive ecosystem of open-source technologies that prepare, process, manage and analyze big data sets [68]. For example, Fig. 6 is the ecosystem of Apache Hadoop [69].

### 08 Mar 2024 v3

# Contemporary Recommendation Systems on Big Data and Their Applications: A Survey

@@ Abstract
This survey paper conducts a comprehensive analysis of the evolution and contemporary landscape of recommendation systems, which have been extensively incorporated across a myriad of web applications. It delves into the progression of personalized recommendation methodologies tailored for online products or services, organizing the array of recommendation techniques into four main categories: content-based, collaborative filtering, knowledge-based, and hybrid approaches, each designed to cater to specific contexts. The document provides an in-depth review of both the historical underpinnings and the cutting-edge innovations in the domain of recommendation systems, with a special focus on implementations leveraging big data analytics. It further outlines and explores the predominant challenges encountered in the current generation of recommendation systems, including issues related to data sparsity, scalability, and the imperative for diversified recommendation outputs. The survey underscores these challenges as promising directions for subsequent research endeavors within the discipline. Additionally, the paper examines various real-life applications driven by recommendation systems, addressing the hurdles involved in seamlessly integrating these systems into everyday life. Ultimately, the survey underscores how the advancements in recommendation systems, propelled by big data technologies, have the potential to significantly enhance real-world experiences.
keywords:
Recommendation System, Big Data, Environment, Management
[…]
[label1]organization=Antai College of Economics and Management, Shanghai Jiao Tong University,addressline=1954 Huashan Road, city=Shanghai, postcode=200030, state=Shanghai, country=China
[…]
[label2]organization=Department of Electrical and Computer Engineering, University of Miami,addressline=1251 Memorial Drive, city=Coral Gables, postcode=33146, state=FL, country=USA
[…]
[label3]organization=Department of Architecture, Cornell University,addressline=129 Sibley Dome, city=Ithaca, postcode=14853, state=NY, country=USA
[…]
[label4]organization=Bascom Palmer Eye Institute, University of Miami Miller School of Medicine,addressline=900 NW 17th St, city=Miami, postcode=33136, state=FL, country=USA
[…]
[label5]organization=Department of Civil & Architectural Engineering, University of Miami,addressline=1251 Memorial Drive, city=Coral Gables, postcode=33146, state=FL, country=USA
[…]
[label6]organization=Department of Architecture, University of Miami,addressline=1223 Theo Dickinson Dr, city=Coral Gables, postcode=33146, state=FL, country=USA
{graphicalabstract}{highlights}
Comprehensive Classification and Analysis of Recommendation Techniques: The paper meticulously categorizes the myriad of recommendation system methodologies into four principal types: content-based, collaborative filtering, knowledge-based, and hybrid approaches. It offers a detailed exploration of how these methodologies have evolved over time and their application in providing personalized recommendations across various online platforms. This classification serves as a foundational framework for understanding the diverse strategies employed in tailoring user experiences and addresses the unique challenges and requirements specific to each approach.
Examination of Challenges and Future Research Directions: Highlighting the current challenges faced by recommendation systems, such as data sparsity, scalability, and the need for diversifying recommendation outputs, the paper not only provides a critical analysis of these issues but also positions them as promising avenues for future research. This focus underscores the dynamic nature of the field and encourages the exploration of innovative solutions to enhance the effectiveness and efficiency of recommendation systems in handling large-scale data and meeting the evolving preferences of users.
Real-World Applications and Impact on Society: The survey extends its analysis beyond theoretical models to examine the practical implementation and impact of recommendation systems in everyday life. It explores their application across various domains including marketing, healthcare, and governance, and discusses how they influence consumer behavior and societal trends. By highlighting the potential of recommendation systems to promote sustainable lifestyles and improve decision-making processes, the paper sheds light on the significant role these technologies play in shaping our digital and real-world experiences.
@@ 1 Introduction
In this survey, we scrutinize the escalating popularity and diverse applications of recommendation systems in web applications, a topic extensively explored by Zhou et al. zhou2021combinatorial . These systems, a specialized category of information filtering systems, are designed to predict user preferences for various items. They play a pivotal role in guiding decision-making processes, such as purchasing decisions and music selections, as discussed by Wang et al. wang2020personalized . An exemplary instance of this application is Amazon’s personalized recommendation engine, which tailors each user’s homepage. Major corporations like Amazon, YouTube, and Netflix utilize these systems to enhance user experience and generate substantial revenue, as highlighted by Adomavicius et al. and Omura et al. adomavicius2005toward ; omura2020ad . Figure 1 from Entezari et al. entezari2021tensor illustrates a modern recommendation system. Additionally, these systems are increasingly pertinent in the field of human-computer interaction (HCI), where they improve interaction efficiency through feedback mechanisms, a topic explored in several studies 7962152 .
Recommendation systems are particularly crucial for certain companies, as their efficiency can lead to substantial revenue generation and competitive advantage, as evidenced in the research by Rismanto et al. and Cui et al. rismanto2020research ; cui2020personalized . For example, Netflix’s “Netflix Prize” challenge aimed to develop a recommender system surpassing their existing algorithm, offering a substantial prize to incentivize innovation.
@@
Furthermore, in the domain of big data, recommendation systems are highly prevalent, as detailed by Li et al. li2021research ; li2021sports . These systems predict user interests in purchasing based on extensive data analysis, including purchase history, ratings, and reviews. There are four widely recognized types of recommendation systems, as identified by Numnonda numnonda2018real : content-based, collaborative filtering-based, knowledge-based, and hybrid-based, each with distinct advantages and drawbacks, as Xiao et al. elucidate xiao2018personalized . For example, collaborative filtering-based systems may face issues such as data sparsity and scalability, as Huang et al. mention huang2019multimodal , and cold-start problems, while content-based systems might struggle to diversify user interests, as noted by Zhang et al. zhang2018mcrs and Benouaret et al. benouaret2020comparative .
As recommendation system algorithms, neural networks, and big data continue to evolve, their integration into various aspects of our lives is increasingly evident. We are witnessing a surge in applications that leverage big data, recommendation systems, and machine learning to enhance our daily experiences. In the realm of marketing, these technologies enable more personalized recommendations that align with individuals’ personal backgrounds and cultural preferences, whether they are watching movies, shopping online, or traveling. The impact extends to governance, where big data applications are facilitating more efficient and responsive public services. Simultaneously, these innovations are promoting a more sustainable lifestyle and fostering stronger connections in healthcare, demonstrating their potential to improve well-being across various sectors. Moreover, these applications are collecting vast amounts of data across all age groups, from children to seniors, empowering individuals with insights to better understand and analyze their behaviors and preferences sun2024said ; huh2019location . This data-driven approach is revolutionizing how we interact with technology, making it a pivotal element in our journey towards a more informed and connected society.
The structure of this paper is outlined as follows: Section 2 provides a comprehensive exploration of recommendation systems, tracing their evolution from historical foundations to contemporary state-of-the-art methodologies, accompanied by a thorough examination of recent advancements within the sector. Section 3 addresses the specific challenges encountered in recommendation systems that leverage big data, such as data sparsity, scalability issues, and the imperative for diverse recommendations. This section also delves into potential solutions for overcoming these obstacles. Section 4 is dedicated to the application of recommendation systems in real-life contexts, discussing their practical implications and the challenges associated with their integration. Finally, the paper culminates with a comprehensive summary presented in Section 7.
@@ 2 Recommendation Systems
Recommendation systems aim to predict users’ preferences for a certain item and provide personalized services yi2019deep . This section will discuss several commonly used recommender methods, such as content-based method, collaborative filtering-based method, knowledge-based method, and hybrid-based method.
@@ 2.1 Content-based Recommendation Systems
The main idea of content-based recommenders is to recommend items based on the similarity between different users or items lops2011content . This algorithm determines and differentiates the main common attributes of a particular user’s favorite items by analyzing the descriptions of those items. Then, these preferences are stored in this user’s profile. The algorithm then recommends items with a higher degree of similarity with the user’s profile. Besides, content-based recommendation systems can capture the specific interests of the user and can recommend rare items that are of little interest to other users. However, since the feature representations of items are designed manually to a certain extent, this method requires a lot of domain knowledge. In addition, content-based recommendation systems can only recommend based on users’ existing interests, so the ability to expand users’ existing interests is limited.
@@ 2.2 Collaborative Filtering-based Recommendation Systems
Collaborative Filtering-based (CF) methods are primarily used in big data processing platforms due to their parallelization characteristics elahi2016survey . The basic principle of the recommendation system based on collaborative filtering is shown in Fig. 2 alhijawi2016using . CF recommendation systems use the behavior of a group of users to recommend to other users al2018similarity . There are mainly two types of collaborative filtering techniques, which are user-based and item-based.
  • 1.
    User-based CF: In the user-based CF recommendation system, users will receive recommendations of products that similar users like rezaimehr2021survey . Many similarity metrics can calculate the similarity between users or items, such as Constrained Pearson Correlation coefficient (CPC), cosine similarity, adjusted cosine similarity, etc. For example, cosine similarity is a measure of similarity between two vectors. Let $x$ and $y$ denote two vectors, cosine similarity between $x$ and $y$ can be represented by
@@
  • 2.
    Item-based CF: Item-based CF algorithm predicts user ratings for items based on item similarity. Generally, item-based CF yields better results than user-based CF because user-based CF suffers from sparsity and scalability issues. However, both user-based CF and item-based CF may suffer from cold-start problems zhang2016fast .
@@
Musto et al. musto2016learning
Volkovs et al.volkovs2017content
Mittal et al.mittal2020smart
Almaguer et al.perez2021content
@@
Zhang et al.zhang2019privacy
Bobadilla et al.bobadilla2020deep
Bobadilla et al.bobadilla2020classification
Rezaimehr et al.rezaimehr2021survey
@@
Dong et al.dong2020interactive
Gazdar et al.gazdar2020new
Alamdari et al.alamdari2020systematic
Cena et al.cena2021logical
@@
Hrnjica et al.hrnjica2020model
Shokeen et al.shokeen2020study
Zagra et al.zagranovskaia2021designing
Ibrahim et al.ibrahim2021hybrid
@@ 2.3 Knowledge-based Recommendation Systems
The main idea of knowledge-based recommendation systems is to recommend items to users based on basic knowledge of users, items, and relationships between items shishehchi2012ontological ; aggarwal2016knowledge . Since knowledge-based recommendation systems do not require user ratings or purchase history, there is no cold start problem for this type of recommendation cabezas2017knowledge . Knowledge-based recommendation systems are well suited for complex domains where items are not frequently purchased, such as cars and apartments tarus2018knowledge . But the acquisition of required domain knowledge can become a bottleneck for this recommendation technique dong2020interactive .
@@ 2.4 Hybrid-based Recommendation Systems
Hybrid-based recommendation systems combine the advantages of multiple recommendation techniques and aim to overcome the potential weaknesses in traditional recommendation systems ribeiro2012pareto . There are seven basic hybrid recommendation techniques ibrahim2021hybrid : weighted, mixed, switching, feature combination, feature augmentation, cascade, and meta-level methods hassan2016enhancing ; zhang2016hybrid . Among all of these methods, the most commonly used is the combination of the CF recommendation methods with other recommendation methods (such as content-based or knowledge-based) to avoid sparsity, scalability, and cold-start problems hrnjica2020model ; zagranovskaia2021designing ; george2019review .
@@ 2.5 Challenges in Modern Recommendation Systems
  • 1.
    Sparsity. As we know, the usage of recommendation systems is growing rapidly. Many commercial recommendation systems use large datasets, and the user-item matrix used for filtering may be very large and sparse. Therefore, the performance of the recommendation process may be degraded due to the cold start problems caused by data sparsity west2016recommendation .
  • 2.
    Scalability. Traditional algorithms will face scalability issues as the number of users and items increases. Assuming there are millions of customers and millions of items, the algorithm’s complexity will be too large. However, recommendation systems must respond to the user’s needs immediately, regardless of the user’s rating history and purchase situation, which requires high scalability. For example, Twitter is a large web company that uses clusters of machines to scale recommendations for its millions of users shokeen2020study .
  • 3.
    Diversity. Recommendation systems also need to increase diversity to help users discover new items. Unfortunately, some traditional algorithms may accidentally do the opposite because they always recommend popular and highly-rated items that some specific users love. Therefore, new hybrid methods need to be developed to improve the performance of the recommendation systems he2021research .
@@ 3 Recommendation System based on Big Data
Big data refers to the massive, high growth rate and diversified information chen2021dqn ; kadam2016big . It requires new processing models to have stronger decision-making and process optimization capabilities acharjya2016survey . Big data has its unique “4V” characteristics, as shown in Fig. 3 zhou2018academic : Volume, Variety, Velocity, and Veracity.
@@ 3.1 Big Data Processing Flow
Big data comes from many sources, and there are many methods to process it ram2018privacy . However, the primary processing of big data can be divided into four steps wu2013data . Besides, Fig. 4 presents the basic flow of big data processing.
  • 1.
@@
  • 2.
    Data Processing and Integration. The collection terminal itself already has a data repository, but it cannot accurately analyze the data. The received information needs to be pre-processed emani2015understandable .
  • 3.
    Data Analysis. In this process, these initial data are always deeply analyzed using cloud computing technology lin2019cloud .
  • 4.
@@ 3.2 Modern Recommendation Systems based on the Big Data
The shortcomings of traditional recommendation systems mainly focus on insufficient scalability and parallelism cheng2020research . For small-scale recommendation tasks, a single desktop computer is sufficient for data mining goals, and many techniques are designed for this type of problems al2021hybrid .
@@
However, the rating data is usually so large for medium-scale recommendation systems that it is impossible to load all the data into memory at once dev2016recommendation . Common solutions are based on parallel computing or collective mining, sampling and aggregating data from different sources, and using parallel computing programming to perform the mining process chen2018disease . The big data processing framework will rely on cluster computers with high-performance computing platforms wan2021research . At the same time, data mining tasks will be deployed on a large number of computing nodes (i.e., clusters) by running some parallel programming tools asiyakeyword , such as MapReduce kadam2016big ; verma2015big . For example, Fig. 5 is the MapReduce in the Recommendation Systems.
In recent years, various big data platforms have emerged uzun2021big . For example, Hadoop and Spark kadam2016big , both developed by the Apache Software Foundation, are widely used open-source frameworks for big data architectures kadam2016big ; zhang2018service . Each framework contains an extensive ecosystem of open-source technologies that prepare, process, manage and analyze big data sets chaithra2015user . For example, Fig. 6 is the ecosystem of Apache Hadoop ait2019distributed .